# A Comprehensive Analysis on LLM-based Node Classification Algorithms

### v1

@@ 4.3 Computational Cost Analysis
We evaluate the training and inference times of various methods in both semi-supervised and supervised settings. Detailed training times are provided in Tables 17 and 18 in the Appendix, while inference times are presented in Table 19. All measurements were conducted on a single NVIDIA H100-80G GPU to ensure consistency.
@@ 5 Fine-grained Analysis Within Each Category
In this section, we present empirical results within each category. For LLM-as-Encoder, we explore the conditions under which LLMs outperform traditional LMs. Additionally, we examine how key components (e.g., model type and size) influence the effectiveness of LLM-as-Reasoner and LLM-as-Predictor.
@@ 5.1 LLM-as-Encoder: Compared with LMs
Motivation and Settings: Both LLMs and small-scale LMs can encode nodes’ associated texts. This raises the question: When do LLMs surpass LMs as encoders? To address this, we evaluate various methods using node features derived from LLMs and LMs, observing the resulting performance differences. For LMs, we select SenBERT-66M (Reimers & Gurevych, 2019) and RoBERTa-355M (Liu et al., 2019). For LLMs, we choose Qwen2.5-3B (Yang et al., 2024) and Mistral-7B (Jiang et al., 2023). The considered methods include: (1) MLP, which solely utilizes node features as input to predict labels without incorporating any graph information, (2) GCN, and (3) GraphSAGE. For each method, we input node features initialized from various LM or LLM backbones while keeping all other components consistent. From the results shown in Table 3 and Table 12 in the Appendix, we can observe that:
Takeaway 8: LLM-as-Encoder significantly outperforms LMs in heterophilic graphs. In heterophilic datasets such as Reddit, LLM-based encoders achieve 2% - 8% higher accuracy than their LM counterparts. This performance gap is most evident with the MLP method (4%–8%). The performance gain is less obvious in homophilic settings. We can also leverage the mutual information (1) to give theoretical insights. In homophilic graphs, edges often connect nodes with the same labels, whereas this does not hold in heterophilic graphs. Therefore, in homophilic graphs, the first term in (1) dominates, and the room for a better encoder, e.g., LLM, to improve is limited.
@@
5.2 LLM-as-Reasoner: Impact of LLM Reasoning Capabilities
Motivation and Settings: In the LLM-as-Reasoner paradigm, as the language model should generate reasoning texts, the adopted language models should be auto-regressive and the model size should be large. To investigate how the advanced reasoning capabilities of LLMs influence overall performance, we replace the default Mistral-7B model in the TAPE method with the more powerful GPT-4o model, keeping all other components unchanged.
Results and Analysis: As presented in Table 4 (semi-supervised settings) and Table 5 (supervised settings), the effectiveness of LLM-as-Reasoner methods positively correlates with the strength of the underlying LLMs. In semi-supervised settings, TAPE utilizing GPT-4o consistently outperforms its Mistral-7B counterpart, achieving performance gains of up to 4% on the Pubmed dataset. However, in supervised scenarios, the performance gap between GPT-4o and Mistral-7B narrows. This reduction is attributed to the abundance of labeled data, which increases the mutual information $I(\mathcal{E};\mathcal{Y}_{l})$ in (1). Consequently, the dependency on node attributes decreases, thereby diminishing the relative advantages of more powerful LLMs.
Based on these findings, we recommend that when abundant supervision is available, practitioners may opt for open-source LLMs instead of more powerful and costlier models for the LLM-as-Reasoner method. This practice can achieve comparable performance without incurring additional costs.
@@ 5.3 LLM-as-Predictor: Sensitivity to LLM Backbones
Motivation and Settings: For most LLM-as-Predictor methods, only open-source LLMs are compatible. Given the diverse choices and varying scales of these models, we aim to investigate the sensitivity of performance to different LLM backbones. This examination seeks to identify potential scaling laws and determine which LLMs excel at the node classification task. Therefore, we choose the best predictor method LLaGA as the baseline, include models of different sizes within the same series, i.e., Qwen2.5-series (Yang et al., 2024). Additionally, we consider similar-scaled models to identify the most suitable for this task, including Qwen2.5-7B, Mistral-7B, and LLaMA3.1-8B. All experiments maintain consistency by only varying the backbone LLMs while keeping other components, training configurations, and hyperparameters unchanged.
Results and Analysis: (1) Scaling within the same series: Comparing Qwen-3B to Qwen-32B (performance shown in Tables 6 and 7, efficiency in Table 14, and performance trends in Figures 3 and 4 in the Appendix), we observe that performance generally improves with larger model sizes. However, beyond Qwen-7B and Qwen-14B, the performance gains become marginal while training and inference times increase significantly. For instance, Qwen-32B takes over 200 milliseconds per sample for inference, which is five times longer than Qwen-7B. Therefore, Qwen-7B or Qwen-14B are recommended as practical choices balancing performance and efficiency. (2) Model selection at similar scales: When comparing models of similar sizes (Tables 6 and 7, and Table 13 in the Appendix), Mistral-7B outperforms other LLMs of comparable scale. Its superior performance makes Mistral-7B the recommended backbone LLM for node classification tasks.
@@ 6 Conclusion
This paper provides guidelines for leveraging LLMs to enhance node classification tasks across diverse real-world applications. We introduce LLMNodeBed, a codebase and testbed for systematic comparisons, featuring ten datasets, eight LLM-based algorithms, and three learning paradigms. Through extensive experiments involving 2,200 models, we uncover key insights: In supervised settings, each category offers unique advantages, but LLM-based approaches deliver marginal improvements over classic methods when ample supervision is available. In zero-shot scenarios, directing powerful LLMs to perform inference with integrated structural context yields the best performance.
Our findings offer practical guidance for practitioners applying LLMs to node classification tasks and highlight research gaps, e.g., the limited exploration of LLMs on heterophilic graphs and the scarcity of such text-rich datasets. We intend to address these gaps in future work. We hope that LLMNodeBed will inspire and serve as a valuable toolkit for further research.
@@
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ A.1 Classic Methods
Early approaches for node classification tasks relied on structural techniques such as Laplacian regularization (Belkin et al., 2006), graph embeddings (Yang et al., 2016a), and label propagation (Zhu et al., 2003). These methods infer node labels by leveraging the connectivity and similarity among nodes within the graph.
@@
Beyond structural methods, node classification can also be approached using LMs by treating each node as a text entity. Fine-tuning LMs (Liu et al., 2019; Reimers & Gurevych, 2019; Wang et al., 2022) allows these models to map textual information directly to node labels, leveraging their strong language understanding capabilities to predict labels. To harness the complementary strengths of GNNs and LMs, i.e., structural and textual information, hybrid LM+GNN architectures have been developed (Jin et al., 2023b; Zhao et al., 2023a; Wen & Fang, 2023). These models integrate LM-encoded textual features with GNN-processed structural features, enhancing node classification performance by combining both modalities.
@@ A.2 LLM-based Methods
LLM as Encoder: LLMs possess an extensive number of parameters, enabling them to generate highly expressive representations. These representations can replace shallow node embeddings, promising to enhance expressiveness and improve the performance of downstream task. A notable approach is ENGINE (Zhu et al., 2024b), which utilizes hidden embeddings from LLMs to construct node embeddings. ENGINE integrates these embeddings with GNNs to propagate and update representations. Specifically, it aggregates hidden embeddings from each LLM layer that processes a node’s text and incorporates them into a cascaded GNN structure.
LLM as Reasoner: A significant advantage of LLMs is their reasoning and planning capabilities (Luo et al., 2024a; Wu et al., 2024). Guided by carefully crafted prompts, LLMs can intelligently execute a variety of downstream tasks. Motivated by this strength, He et al. (2023) introduced TAPE, a method that leverages LLM as Reasoner for node classification. Specifically, TAPE prompts the LLM to generate predictions along with a chain-of-thought reasoning process that includes explanations, denoted as $s_{v}^{\text{exp}}=\text{LLM}(s_{v}^{\text{orig}},p)$, where $p$ denotes the textual prompt, and $s_{v}^{\text{orig}}$ and $s_{v}^{\text{exp}}$ represent the original and generated texts for node $v$, respectively. Both the original and generated texts are processed by an LM to produce embeddings as $\bm{x}_{v}^{\text{orig}}=\text{LM}(s_{v}^{\text{orig}})$ and $\bm{x}_{v}^{\text{exp}}=\text{LM}(s_{v}^{\text{exp}})$, which are subsequently processed by GNNs for the classification task as:
@@
Subsequent works have enhanced the reliability of LLM reasoning. For example, KEA (Chen et al., 2024b) prompts LLMs to extract and explain specific technical terms from a node’s original text instead of making direct predictions, thereby mitigating potential misguidance. Overall, the LLM-as-Reasoner paradigm leverages LLMs’ reasoning capabilities to generate reliable explanations, thus augmenting the original graph data.
@@
In the LLaGA framework, only the parameters of the projection layer are tuned, utilizing the next-token-prediction loss based on ground-truth labels and generated outputs. GraphGPT employs a more complex framework with three distinct pre-training and instruction tuning stages. Other LLM-as-Predictor methods (Chai et al., 2023; Perozzi et al., 2024; Kong et al., 2024; Huang et al., 2024; Zhao et al., 2023b; Ji et al., 2024) share similar frameworks with LLaGA but vary in integration approaches, training objectives, and tackled tasks.

### v2

@@ 4.3 Computational Cost Analysis
Time Cost: We evaluate the training and inference times of various methods in supervised settings. Detailed training times are provided in Table 19 in the Appendix, while inference times are presented in Table 20. All measurements were conducted on a single NVIDIA H100-80G GPU to ensure consistency.
@@
Memory Cost: We also analyze the memory costs of different methods during both the training and inference stages to assess their resource usage. We select three datasets, Cora, WikiCS, and arXiv, spanning a range from 2,708 to 169,343 nodes. The results, presented in Table 21 in the Appendix, were measured using a single H100-80GB GPU to ensure consistency and comparability. The results show that LLM-as-Encoder methods demonstrate memory costs comparable to traditional GNNs, making them practical choices for deployment. In contrast, methods involving fine-tuning an LM or LLM-as-Predictor methods are significantly more memory-intensive, largely due to the high parameterization of language models. These findings underscore the importance of addressing efficiency challenges to enable the practical deployment of LLM-based methods.
@@ 5 Fine-grained Analysis Within Each Category
In this section, we present empirical results within each category. For LLM-as-Encoder, we explore the conditions under which LLMs outperform traditional LMs. Additionally, we examine how key components (e.g., model type and size) influence the effectiveness of LLM-as-Explainer and LLM-as-Predictor.
@@ 5.1 LLM-as-Encoder: Compared with LMs
[…]
Semi-supervised Supervised
Method Encoder Cornell Texas Wisconsin Washington Cornell Texas Wisconsin Washington
Homophily Ratio (%) 11.55 6.69 16.27 17.07 11.55 6.69 16.27 17.07
MLP SenBERT 50.59±${}_{\pm\text{3.14}}$ 56.67±${}_{\pm\text{2.15}}$ 71.98±${}_{\pm\text{1.59}}$ 63.26±${}_{\pm\text{2.89}}$ 66.15±${}_{\pm\text{1.92}}$ 76.32±${}_{\pm\text{3.72}}$ 81.51±${}_{\pm\text{7.00}}$ 70.44±${}_{\pm\text{8.65}}$
RoBERTa 59.08±${}_{\pm\text{2.57}}$ 67.47±${}_{\pm\text{1.29}}$ 73.87±${}_{\pm\text{1.62}}$ 65.43±${}_{\pm\text{3.44}}$ 66.67±${}_{\pm\text{8.88}}$ 74.21±${}_{\pm\text{6.09}}$ 80.00±${}_{\pm\text{9.88}}$ 76.96±${}_{\pm\text{7.48}}$
Qwen-3B 57.78±${}_{\pm\text{3.24}}$ 76.27±${}_{\pm\text{1.61}}$ 82.36±${}_{\pm\text{1.62}}$ 75.11±${}_{\pm\text{1.92}}$ 77.95±${}_{\pm\text{4.76}}$ 88.95±${}_{\pm\text{3.07}}$ 88.68±${}_{\pm\text{6.64}}$ 83.48±${}_{\pm\text{1.74}}$
Mistral-7B 59.87±${}_{\pm\text{6.72}}$ 76.27±${}_{\pm\text{1.08}}$ 83.30±${}_{\pm\text{1.42}}$ 74.24±${}_{\pm\text{0.88}}$ 78.46±${}_{\pm\text{4.17}}$ 90.53±${}_{\pm\text{3.16}}$ 89.43±${}_{\pm\text{5.15}}$ 83.91±${}_{\pm\text{5.60}}$
GCN SenBERT 46.80±${}_{\pm\text{2.13}}$ 54.93±${}_{\pm\text{0.68}}$ 58.30±${}_{\pm\text{2.56}}$ 52.61±${}_{\pm\text{1.35}}$ 50.77±${}_{\pm\text{10.18}}$ 59.47±${}_{\pm\text{5.16}}$ 61.13±${}_{\pm\text{8.65}}$ 61.30±${}_{\pm\text{1.62}}$
RoBERTa 47.06±${}_{\pm\text{2.19}}$ 55.20±${}_{\pm\text{2.78}}$ 54.91±${}_{\pm\text{3.40}}$ 54.89±${}_{\pm\text{1.50}}$ 51.79±${}_{\pm\text{7.68}}$ 58.42±${}_{\pm\text{7.33}}$ 59.24±${}_{\pm\text{8.82}}$ 61.31±${}_{\pm\text{5.39}}$
Qwen-3B 53.59±${}_{\pm\text{2.07}}$ 56.80±${}_{\pm\text{4.29}}$ 63.02±${}_{\pm\text{2.16}}$ 64.56±${}_{\pm\text{4.06}}$ 58.46±${}_{\pm\text{10.56}}$ 64.74±${}_{\pm\text{7.37}}$ 65.28±${}_{\pm\text{6.82}}$ 67.83±${}_{\pm\text{3.74}}$
Mistral-7B 54.64±${}_{\pm\text{1.52}}$ 58.67±${}_{\pm\text{3.60}}$ 62.08±${}_{\pm\text{2.61}}$ 61.52±${}_{\pm\text{3.61}}$ 59.49±${}_{\pm\text{6.96}}$ 65.79±${}_{\pm\text{6.66}}$ 64.90±${}_{\pm\text{5.67}}$ 66.96±${}_{\pm\text{4.84}}$
SAGE SenBERT 52.55±${}_{\pm\text{1.58}}$ 61.73±${}_{\pm\text{1.37}}$ 70.47±${}_{\pm\text{1.75}}$ 65.54±${}_{\pm\text{2.44}}$ 68.72±${}_{\pm\text{4.97}}$ 80.00±${}_{\pm\text{5.91}}$ 83.02±${}_{\pm\text{6.31}}$ 76.96±${}_{\pm\text{4.88}}$
RoBERTa 55.55±${}_{\pm\text{3.44}}$ 64.26±${}_{\pm\text{6.26}}$ 73.59±${}_{\pm\text{2.72}}$ 66.08±${}_{\pm\text{1.60}}$ 70.26±${}_{\pm\text{8.37}}$ 80.53±${}_{\pm\text{2.68}}$ 81.89±${}_{\pm\text{7.42}}$ 74.35±${}_{\pm\text{7.95}}$
Qwen-3B 57.13±${}_{\pm\text{2.29}}$ 78.53±${}_{\pm\text{1.76}}$ 83.21±${}_{\pm\text{1.39}}$ 72.18±${}_{\pm\text{3.66}}$ 74.87±${}_{\pm\text{2.99}}$ 89.47±${}_{\pm\text{1.67}}$ 91.32±${}_{\pm\text{2.82}}$ 83.48±${}_{\pm\text{3.25}}$
Mistral-7B 56.86±${}_{\pm\text{1.37}}$ 76.53±${}_{\pm\text{2.40}}$ 83.96±${}_{\pm\text{1.55}}$ 73.91±${}_{\pm\text{0.97}}$ 77.44±${}_{\pm\text{2.99}}$ 91.05±${}_{\pm\text{2.69}}$ 89.44±${}_{\pm\text{4.24}}$ 81.74±${}_{\pm\text{4.48}}$
H2GCN SenBERT 56.34±${}_{\pm\text{1.67}}$ 66.67±${}_{\pm\text{2.95}}$ 73.40±${}_{\pm\text{1.68}}$ 70.55±${}_{\pm\text{4.95}}$ 73.85±${}_{\pm\text{7.14}}$ 84.21±${}_{\pm\text{4.40}}$ 86.42±${}_{\pm\text{6.01}}$ 77.83±${}_{\pm\text{7.20}}$
RoBERTa 60.00±${}_{\pm\text{3.54}}$ 68.13±${}_{\pm\text{2.93}}$ 75.66±${}_{\pm\text{2.12}}$ 71.52±${}_{\pm\text{1.22}}$ 74.87±${}_{\pm\text{7.68}}$ 83.16±${}_{\pm\text{6.14}}$ 84.53±${}_{\pm\text{9.04}}$ 79.13±${}_{\pm\text{5.43}}$
Qwen-3B 61.57±${}_{\pm\text{3.89}}$ 80.13±${}_{\pm\text{6.45}}$ 84.53±${}_{\pm\text{0.70}}$ 74.67±${}_{\pm\text{1.77}}$ 76.41±${}_{\pm\text{2.99}}$ 92.11±${}_{\pm\text{2.88}}$ 89.81±${}_{\pm\text{3.29}}$ 85.22±${}_{\pm\text{3.99}}$
Mistral-7B 59.22±${}_{\pm\text{4.54}}$ 72.93±${}_{\pm\text{8.21}}$ 81.89±${}_{\pm\text{1.51}}$ 68.59±${}_{\pm\text{4.46}}$ 75.89±${}_{\pm\text{3.84}}$ 89.47±${}_{\pm\text{3.72}}$ 89.43±${}_{\pm\text{5.42}}$ 86.09±${}_{\pm\text{3.25}}$
@@
Motivation and Settings: Both LLMs and small-scale LMs can encode nodes’ associated texts. This raises the question: When do LLMs surpass LMs as encoders? To address this, we evaluate various methods using node features derived from LLMs and LMs, observing the resulting performance differences. For LMs, we select SenBERT-66M (Reimers & Gurevych, 2019) and RoBERTa-355M (Liu et al., 2019). For LLMs, we choose Qwen2.5-3B (Qwen et al., 2025) and Mistral-7B (Jiang et al., 2023). The evaluated datasets include both homophilic and heterophilic graphs to ensure generality. For heterophilic graphs, we use the released datasets collected by Wang et al. (2025) [^1] https://huggingface.co/datasets/Graph-COM/Text-Attributed-Graphs with processing details explained in Appendix C.1. The considered methods include: (1) MLP, which solely utilizes node features as input to predict labels without incorporating any graph information, (2) GCN, (3) GraphSAGE, and (4) H2GCN (Zhu et al., 2020), a representative GNN explicitly designed for heterophilic settings, included specifically to accommodate heterophilic graphs. For each method, we initialize node features using embeddings derived from various LM or LLM backbones, ensuring that all other components remain consistent across evaluations.
From the results shown in Table 4 (homophilic graphs) and Table 5 (heterophilic graphs), we can observe that: Takeaway 8: LLM-as-Encoder significantly outperforms LMs in less informative graphs, e.g., heterophilic ones.
For homophilic graphs, the performance gap between LLMs and LMs is relatively small and becomes noticeable only in graphs with lower homophily, such as WikiCS and Reddit. For instance, on Reddit, LLM-based encoders achieve 4% higher accuracy than their LM counterparts. The performance gap is substantially larger for heterophilic graphs, with differences reaching up to 10% across methods in both semi-supervised and supervised settings. For example, on the Texas dataset, LLM-based encoders achieve a 12% improvement over their LM counterparts for the GCN, SAGE, and H2GCN methods in the semi-supervised setting. We can also explain these results using mutual information in Equation (1). In homophilic graphs, edges often connect nodes with the same labels. Consequently, the first term in (1) dominates, leaving limited room for improved encoders like LLMs to enhance performance. In contrast, this property does not hold for less-informative heterophilic graphs, where neighboring nodes are less likely to share the same labels, making the choice of encoder significantly more impactful.
5.2 LLM-as-Explainer: Impact of LLM Capabilities
Motivation and Settings: In the LLM-as-Explainer paradigm, as the language model should generate explanatory texts, the adopted language models should be auto-regressive and the model size should be large. To investigate how the advanced reasoning capabilities of LLMs influence overall performance, we replace the default Mistral-7B model in the TAPE method with the more powerful GPT-4o model, keeping all other components unchanged.
@@ 5.2 LLM-as-Explainer: Impact of LLM Capabilities
Results and Analysis: As presented in Table 6 (semi-supervised settings) and Table 7 (supervised settings), the effectiveness of LLM-as-Explainer methods positively correlates with the strength of the underlying LLMs. In semi-supervised settings, TAPE utilizing GPT-4o consistently outperforms its Mistral-7B counterpart, achieving performance gains of up to 4% on the Pubmed dataset. However, in supervised scenarios, the performance gap between GPT-4o and Mistral-7B narrows. This reduction is attributed to the abundance of labeled data, which increases the mutual information $I(\mathcal{E};\mathcal{Y}_{l})$ in (1). Consequently, the dependency on node attributes decreases, thereby diminishing the relative advantages of more powerful LLMs.
Based on these findings, we recommend that when abundant supervision is available, practitioners may opt for open-source LLMs instead of more powerful and costlier models for the LLM-as-Explainer method. This practice can achieve comparable performance without incurring additional costs.
@@ 5.3 LLM-as-Predictor: Sensitivity to LLM Backbones
Motivation and Settings: For most LLM-as-Predictor methods, only open-source LLMs are compatible. Given the diverse choices and varying scales of these models, we aim to investigate the sensitivity of performance to different LLM backbones. This examination seeks to identify potential scaling laws and determine which LLMs excel at the node classification task. Therefore, we choose the best predictor method LLaGA as the baseline, include models of different sizes within the same series, i.e., Qwen2.5-series (Qwen et al., 2025). Additionally, we consider similar-scaled models to identify the most suitable for this task, including Qwen2.5-7B, Mistral-7B, and LLaMA3.1-8B. All experiments maintain consistency by only varying the backbone LLMs while keeping other components, training configurations, and hyperparameters unchanged.
Results and Analysis: (1) Scaling within the same series: Comparing Qwen-3B to Qwen-32B (performance shown in Tables 8 and 9, efficiency in Table 16, and performance trends in Figures 3 and 4 in the Appendix), we observe that performance generally improves with larger model sizes. However, beyond Qwen-7B and Qwen-14B, the performance gains become marginal while training and inference times increase significantly. For instance, Qwen-32B takes over 200 milliseconds per sample for inference, which is five times longer than Qwen-7B. Therefore, Qwen-7B or Qwen-14B are recommended as practical choices balancing performance and efficiency. (2) Model selection at similar scales: When comparing models of similar sizes (Tables 8 and 9, and Table 15 in the Appendix), Mistral-7B outperforms other LLMs of comparable scale. Its superior performance makes Mistral-7B the recommended backbone LLM for node classification tasks.
@@ 6 Conclusion
This paper provides guidelines for leveraging LLMs to enhance node classification tasks across diverse real-world applications. We introduce LLMNodeBed, a codebase and testbed for systematic comparisons, featuring 14 datasets, 8 LLM-based algorithms, 8 classic algorithms, and 3 learning paradigms. Through extensive experiments involving 2,700 models, we uncover key insights: In supervised settings, each category offers unique advantages, but LLM-based approaches deliver marginal improvements over classic methods when ample supervision is available. In zero-shot scenarios, directing powerful LLMs to perform inference with integrated structural context yields the best performance.
Our findings offer practical guidance for practitioners applying LLMs to node classification tasks and highlight research gaps, e.g., the limited exploration of LLMs on heterophilic graphs and the scarcity of such text-rich datasets. We hope that LLMNodeBed will inspire and serve as a valuable toolkit for further research.
Acknowledgments
This research is supported in part by project #MMT-p2-23 of the Shun Hing Institute of Advanced Engineering, The Chinese University of Hong Kong, by grants from the Research Grants Council of the Hong Kong SAR, China (No. CUHK 14217622). The authors would like to express their gratitude to the reviewers for their feedback, which has improved the clarity and contribution of the paper.
@@ Acknowledgments
In this paper, we did not use any non-public data, unauthorized software, or APIs, and there are no privacy or other related ethical concerns associated with our work. Similar to other machine learning models designed for node classification, our LLM-based algorithms have the potential to be misused for malicious purposes, such as unauthorized surveillance, manipulation of social networks, or exploitation of financial systems. We pledge to restrict the usage of our models exclusively to research settings to prevent such misuse and encourage the responsible development and deployment of these technologies in real-world applications.
@@ A.1 Classic Methods
Early approaches for node classification tasks relied on structural techniques such as Laplacian regularization (Belkin et al., 2006), graph embeddings (Yang et al., 2016), and label propagation (Zhu et al., 2003). These methods infer node labels by leveraging the connectivity and similarity among nodes within the graph.
@@
Beyond structural methods, node classification can also be approached using LMs by treating each node as a text entity. Fine-tuning LMs (Liu et al., 2019; Reimers & Gurevych, 2019; Wang et al., 2022) allows these models to map textual information directly to node labels, leveraging their strong language understanding capabilities to predict labels. To harness the complementary strengths of GNNs and LMs, i.e., structural and textual information, hybrid LM+GNN architectures have been developed (Jin et al., 2023b; Zhao et al., 2023a; Wen & Fang, 2023; Li et al., 2023a). These models integrate LM-encoded textual features with GNN-processed structural features, enhancing node classification performance by combining both modalities.
@@ A.2 LLM-based Methods
LLM as Encoder: LLMs possess an extensive number of parameters, enabling them to generate highly expressive representations. These representations can replace shallow node embeddings, promising to enhance expressiveness and improve the performance of downstream task. A notable approach is ENGINE (Zhu et al., 2024), which utilizes hidden embeddings from LLMs to construct node embeddings. ENGINE integrates these embeddings with GNNs to propagate and update representations. Specifically, it aggregates hidden embeddings from each LLM layer that processes a node’s text and incorporates them into a cascaded GNN structure.
LLM as Explainer: A notable advantage of LLMs lies in their generative capabilities, enabling them to intelligently perform a wide range of downstream tasks, from creative generation to reasoning and planning (Luo et al., 2024a; Wu et al., 2024). Motivated by this strength, He et al. (2024) introduced TAPE, a method that leverages LLM as Explainer for node classification. Specifically, TAPE prompts the LLM to generate predictions along with a chain-of-thought reasoning process that includes explanations, denoted as $s_{v}^{\text{exp}}=\text{LLM}(s_{v}^{\text{orig}},p)$, where $p$ denotes the textual prompt, and $s_{v}^{\text{orig}}$ and $s_{v}^{\text{exp}}$ represent the original and generated texts for node $v$, respectively. Both the original and generated texts are processed by an LM to produce embeddings as $\bm{x}_{v}^{\text{orig}}=\text{LM}(s_{v}^{\text{orig}})$ and $\bm{x}_{v}^{\text{exp}}=\text{LM}(s_{v}^{\text{exp}})$, which are subsequently processed by GNNs for the classification task as:
@@
Subsequent works have enhanced the reliability of LLM-generated explanations (Fang et al., 2024). For example, KEA (Chen et al., 2024b) prompts LLMs to extract and explain specific technical terms from a node’s original text instead of making direct predictions, thereby mitigating potential misguidance. Overall, the LLM-as-Explainer paradigm harnesses the reasoning and generative capabilities of LLMs to produce reliable explanations, thereby enriching the original graph data and enhancing downstream tasks.
@@
In the LLaGA framework, only the parameters of the projection layer are tuned, utilizing the next-token-prediction loss based on ground-truth labels and generated outputs. GraphGPT employs a more complex framework with three distinct pre-training and instruction tuning stages. Other LLM-as-Predictor methods (Chai et al., 2023; Perozzi et al., 2024; Kong et al., 2025; Huang et al., 2024; Zhao et al., 2023b; Ji et al., 2024) share similar frameworks with LLaGA but vary in integration approaches, training objectives, and tackled tasks.